Paul Swanson has an opportunity to acquire a franchise from The Yogurt Place, Inc., to dispense frozen yogurt products under The Yogurt Place name. Mr. Swanson has assembled the following information relating to the franchise:

  1. A suitable location in a large shopping mall can be rented for $3,100 per month.
  2. Remodeling and necessary equipment would cost $294,000. The equipment would have a 20-year life and a $14,700 salvage value. Straight-line depreciation would be used, and the salvage value would be considered in computing depreciation.
  3. Based on similar outlets elsewhere, Mr. Swanson estimates that sales would total $340,000 per year. Ingredients would cost 20% of sales.
  4. Operating costs would include $74,000 per year for salaries, $3,900 per year for insurance, and $31,000 per year for utilities. In addition, Mr. Swanson would have to pay a commission to The Yogurt Place, Inc., of 14.5% of sales.

Required:

1. Prepare a contribution format income statement that shows the expected net operating income each year from the franchise outlet.

2-a. Compute the simple rate of return promised by the outlet.

2-b. If Mr. Swanson requires a simple rate of return of at least 18%, should he acquire the franchise?

3-a. Compute the payback period on the outlet.

3-b. If Mr. Swanson wants a payback of three years or less, will he acquire the franchise?
Answer #1

SOLUTION

1.

Sales 340,000
Variable expenses:
Cost of ingredients (20%* 340,000) 68,000
Commissions (14.5%*340,000) 49,300 117,300
Contribution Margin 222,700
Selling and administrative expenses-
Salaries 74,000
Depreciation(294,000-14,700)/20 13,965
Insurance 3,900
Utilities 31,000
Rent (3,100*12) 37,200 160,065
Net Operating Income 62,635

2A) Simple rate of return = Net operating income / Initial Investment

= 62,635 / 294,000

= 21.30%

2B) As the rate of return is higher than the required 18%, he should take the franchise.

3A) Payback period = Initial Investment/ Annual net cash flow

= 294,000 / 76,600 = 3.8 years

Annual net cash flow = Net operating income + Depreciation

= 62,635 + 13,965 = 76,600

3B) As the payback period is more than 3 years, he will not acquire the franchise.